OpenStudy (anonymous):

Plz help me ill give u medals triangle RST is an equilateral triangle; . line TX is perpendicular of line RS. If RT = 8 , then TX= ? a. 3 sqrt(3) b. 4 sqrt(3) c. sqrt(5) d. 3 sqrt(2)

OpenStudy (anonymous):

this is what i came up with: 9.2378752886836027713625866050808

OpenStudy (anonymous):

anser is b

OpenStudy (anonymous):

Equilateral triangles are also isosceles, and if a sgmt from a vertex is perpendicular to the base, then the base is bisected. Refer to picture. Then we can use pythagorean's thm on the right triangle with leg 4 and hyp 8 to solve for TX. 4² + TX² = 8² TX² = 64 - 16 TX² = 48 TX = sqrt(16*3) TX = 4sqrt(3)

OpenStudy (anonymous):